# HyperEvent: Learning Cohesive Events for Large-scale Dynamic Link Prediction

### v1

@@ 3.3 HyperEvent Framework
Figure 2 presents the methodological framework of the HyperEvent. The framework consists of four steps:1) Real-time Adjacency Table Maintenance; 2) Hyper-Event Extraction; 3) Pairwise Correlation Encoding; 4) Authenticity Discrimination. Subsequent sections detail the design of each component.
Real-time Adjacency Table Maintenance (circled-1 in Figure 2)
Inspired by the embedding updates in node-centric methods (Section 3.2), HyperEvent maintains a dynamic adjacency table $\mathcal{A}$. For each node $v\in\mathcal{V}$, $\mathcal{A}_{v}$ stores the $n_{\text{neighbor}}$ most recent interaction partners ordered chronologically:
@@
where $t>\tau_{1}>\tau_{2}>\cdots>\tau_{k}$ and $k\leq n_{\text{neighbor}}$. Here, $n_{\text{neighbor}}$ controls the storage overhead of the adjacency table and critically defines the historical depth considered when constructing hyper-events for a node, enabling efficient aggregation of relevant events. Specifically, a smaller $n_{\text{neighbor}}$ value prioritizes short-term historical interactions, enhancing sensitivity to immediate dynamics, whereas a larger value incorporates long-term contextual patterns, capturing enduring relationships. Practical event prediction scenarios exhibit distinct preferences: for instance, in highly dynamic environments like social networks where interactions evolve rapidly, a lower $n_{\text{neighbor}}$ (e.g., near 10) is preferable to adapt swiftly to recent changes; conversely, in stable relational settings such as collaboration networks, a higher $n_{\text{neighbor}}$ (e.g., near 50) better utilizes extensive historical data to improve accuracy. Empirical results shows that $n_{\text{neighbor}}=$10-50 suffices for most datasets, balancing computational efficiency and contextual richness.
@@
Hyper-event Extraction (circled-2 in Figure 2)
@@
The hyper-event candidate $\mathcal{H}_{e^{*}}$ is constructed as:
@@
Figure 3: Event correlation vector.
Pairwise Correlation Encoding (circled-3 in Figure 2)
As shown in figure 3, for each event $e=(u,v,t)\in\mathcal{H}_{e^{*}}$, compute a four-dimensional relational vector $\mathbf{r}$ capturing its structural-temporal correlations with $e^{*}$:
$\mathbf{r}=\big{[}d(u^{*},u),\,d(u^{*},v),\,d(v^{*},u),\,d(v^{*},v)\big{]}^{\top}.$ (12)
$d(a,b)$ is a distance metric that measures the overlap of 1-hop temporal neighborhoods between two nodes.
$d(a,b)=\frac{1}{|\mathcal{A}_{a}|\cdot|\mathcal{A}_{b}|}\sum_{i=1}^{|\mathcal{A}_{a}|}\sum_{j=1}^{|\mathcal{A}_{b}|}\mathbb{I}\big{[}\mathcal{A}_{a}(i)=\mathcal{A}_{b}(j)\big{]},$ (13)
where $\mathbb{I}$ is the indicator function. This yields a $4$-dimensional relational feature sequence $\mathbf{R}_{\mathcal{H}_{e^{*}}}=[\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{|\mathcal{H}_{e^{*}}|}]$
Enhanced Correlation Vector Sequence: To comprehensively capture the higher-order correlations, we further propose 0-hop and 2-hop correlation.
i) 0-hop correlation: Measures the direct co-occurrence frequency between nodes by computing the proportion of a node’s temporal neighbors that equal the target node. Formally:
@@
This quantifies asymmetric dependence, where $d^{(0)}(a,b)$ reflects the likelihood of $a$ appearing in $b$’s direct historical interactions.
ii) 2-hop correlation: Evaluates 2-hop neighborhood similarities while maintaining computational tractability. The neighborhood expansion is constrained to top $\lfloor\sqrt{n_{\text{neighbor}}}\rfloor$ neighbors $\tilde{\mathcal{A}}_{a}$, then compute:
@@
with $\tilde{\mathcal{A}}^{(2)}_{a}=\bigcup_{x\in\tilde{\mathcal{A}}_{a}}\tilde{\mathcal{A}}_{x}$. $\tilde{\mathcal{A}}$ means top-$\lfloor\sqrt{n_{\text{neighbor}}}\rfloor$ part of $\mathcal{A}$. The $\sqrt{\cdot}$ ensures computational complexity aligns with the 1-hop metric.
The integrated relational vector concatenates correlations for all node pairs between $e$ and $e^{*}$:
$\mathbf{r}=\begin{aligned} \end{aligned}$ (16)
This $12$-dimensional representation holistically encodes multi-scale structural dependencies within the hyper-event candidate. The relational sequence $\mathbf{R}_{\mathcal{H}_{e^{*}}}=[\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{|\mathcal{H}_{e^{*}}|}]$ preserves chronological order of events in $\mathcal{H}_{e^{*}}$, enabling downstream models to exploit temporal dynamics while mitigating information loss in higher-order neighborhood abstraction.
Although the complexity of Equations (13) and (15) is $O(n_{\text{neighbor}}^{2})$, this remains efficient due to GPU-accelerated tensor parallelism and the small, empirically configured neighborhood size $n_{\text{neighbor}}\in[10,50]$.
Figure 4: Parallelized training framework. Adjacency tables precomputed at key timestamps ($t_{1},t_{2},t_{3}$).
Hyper-event Discriminator
Finally, an Transformer-based encoder is used as the hyper-event discriminator , which processes $\mathbf{R}_{\mathcal{H}_{e^{*}}}$ to classify $\mathcal{H}_{e^{*}}$’s authenticity:
$D(\mathcal{H}_{e^{*}})=\sigma\big{(}\mathbf{w}^{\top}\text{Transformer}(\mathbf{R}_{\mathcal{H}_{e^{*}}})+b\big{)},$ (17)
where $\sigma$ denotes the sigmoid function. Binary cross-entropy loss is employed to optimize the discriminator:
@@
where $y=1$ if the query event $e^{*}$ represents a true interaction, and $y=0$ otherwise.
Advantages over Event-centric Methods: 1). Feature Agnosticism: Eliminates dependency on pre-defined features by deriving relational vectors purely from interaction history. 2). Holistic Pattern Modeling: Explicitly encodes inter-event dependencies within $\mathcal{H}_{e^{*}}$ via $\mathbf{r}_{k}$, overcoming atomic aggregation. 3). Relational Inductivity: Temporal cross-distance $d(\cdot,\cdot)$ intrinsically models structural-temporal relationships. 4). Constant-Time Extraction: $\mathcal{O}(1)$ hyper-event retrieval via $\mathcal{A}$ replaces $\mathcal{O}(|\mathcal{E}|)$ neighborhood sampling. 5). Scalability: Enables graph segmentation and parallel training through pre-calculated adjacency tables.
3.4 Efficient Training Algorithm
@@
The HyperEvent architecture fundamentally differs from existing methods through its dependence on adjacency tables rather than learned node states. This enables segmentation-based training: given an adjacency table at any starting timestamp, HyperEvent can commence training from that precise temporal point without loss of continuity. Consequently, precomputed adjacency tables are only required at segment boundaries, which are efficiently obtainable through cumulative aggregation.
@@
Finally, the event streams in each segment can be trained in parallel. As illustrated in Figure 4, the proposed method partitions the original event stream $\mathcal{G}$ into multiple non-overlapping segments. Each segment corresponds to a distinct time interval demarcated by split points. Crucially, adjacency tables are precomputed at the starting timestamp of every segment, capturing the cumulative graph state up to that point. This precomputation mechanism, detailed in Section 3.3, ensures that the cumulative structural information captured at segmentation boundaries is functionally equivalent to that obtained during continuous sequential processing.
@@
Efficient Implementation. HyperEvent achieves exceptional computational efficiency, which ensures scalability and real-time applicability in large-scale temporal graphs. The exhaustive complexity analyses are elaborated in the supplementary material. Furthermore, the parallel training paradigm fundamentally leverages tensorization rather than conventional multi-threading implementation. Crucially, all computations within the framework operate directly on tensors, inherently enabling batched parallel execution on GPUs. Specifically, the original event stream tensor, shaped as ($|\mathcal{E}|$, 2), is partitioned into segments, transforming it into a ($n_{\text{segment}}$, $\lceil|\mathcal{E}|/n_{\text{segment}}\rceil$, 2) tensor, where padding with a designated value (-1) is applied to handle event counts not perfectly divisible by the segment count. Computational operations defined for HyperEvent naturally extend by introducing an additional segment dimension. This tensor-centric parallelism bypasses the computational overhead intrinsic to CPU multi-threading and eliminates constraints imposed by finite CPU core counts.
4 Experiment
4.1 Experimental Setup
HyperEvent is evaluated using the Temporal Graph Benchmark (TGB) (Huang et al. 2023), adhering to its standardized evaluation protocols with filtered Mean Reciprocal Rank (MRR) as the metric. TGB provides five benchmark datasets for dynamic link prediction: tgbl-wiki-v2 (Wiki), tgbl-review-v2 (Review), tgbl-coin-v2 (Coin), tgbl-comment (Comment), and tgbl-flight-v2 (Flight). Notably, Comment ( 44 million events) and Flight ( 67 million events) represent large-scale graphs. Optimal hyperparameters were determined via grid search. As the Transformer architecture within HyperEvent is not a contribution of this work, a fixed configuration was employed. Further details regarding datasets, evaluation metrics, specific hyperparameters, training, and experimental hardware are provided in the supplementary material. For clarity in tables and figures, all reported MRR values are presented as percentages (i.e., scaled by 100).
4.2 Baselines
HyperEvent is compared against the following temporal graph representation learning and dynamic link prediction methods, as introduced in Related Works:
[…]
Deep learning methods: JODIE (Kumar, Zhang, and Leskovec 2019), DyRep (Rakshit Trivedi and Zha 2019), TGAT (Xu et al. 2020), TGN (Rossi et al. 2020), CAWN (Wang et al. 2021b), NAT (Luo and Li 2022), TCL (Wang et al. 2021a), GraphMixer (Cong et al. 2023), DyGFormer (Yu et al. 2023), TNCN (Zhang et al. 2024), CTAN (Gravina et al. 2024).
Model Wiki Review Coin Comment Flight
EdgeBank 49.5 2.3 45.2 12.9 16.7
[…]
JODIE 63.1$\pm$1.7 34.7$\pm$0.1 OOM OOM OOM
DyRep 5.0$\pm$1.7 22.0$\pm$3.0 45.2$\pm$4.6 28.9$\pm$3.3 55.6$\pm$1.4
TGAT 14.1$\pm$0.7 35.5$\pm$1.2 OOM OOM OOM
TGN 39.6$\pm$6.0 34.9$\pm$2.0 58.6$\pm$3.7 37.9$\pm$2.1 70.5$\pm$2.0
CAWN 71.1$\pm$0.6 19.3$\pm$0.1 OOM OOM OOM
NAT 74.9$\pm$1.0 34.1$\pm$2.0 OOM OOM OOM
TCL 20.7$\pm$2.5 19.3$\pm$0.9 OOM OOM OOM
GraphMixer 11.8$\pm$0.2 52.1$\pm$1.5 OOM OOM OOM
DyGFormer 79.8$\pm$0.4 22.4$\pm$1.5 75.2$\pm$0.4 67.0$\pm$0.1 OOM
TNCN 71.8$\pm$0.4 37.7$\pm$1.0 76.2$\pm$0.4 69.7$\pm$0.6 82.0$\pm$0.4
CTAN 66.8$\pm$0.7 40.5$\pm$0.4 74.8$\pm$0.4 67.1$\pm$6.7 OOM
HyperEvent 80.9$\pm$0.1 26.8$\pm$0.4 77.3$\pm$0.3 75.9$\pm$0.2 87.7$\pm$0.3
Table 1: MRR of methods on different datasets. The best-performing method on each dataset is highlighted in boldface, while the second-best is indicated with an underline. ‘OOM’ denotes out-of-memory errors.
4.3 Baseline Comparison
Table 1 presents the MRR of HyperEvent against 12 state-of-the-art baselines. HyperEvent achieves state-of-the-art performance on four datasets: Wiki (80.9), Coin (77.3), Comment (75.9), and Flight (87.7). Crucially, it surpasses the current top methods on the TGB leaderboard by significant margins [^1] https://tgb.complexdatalab.com/docs/leader˙linkprop/, improving MRR by 1.1% (vs. DyGFormer) on Wiki, 1.1% (vs. TNCN) on Coin, 6.2% (vs. TNCN) on Comment, and 5.7% (vs. TNCN) on Flight. These gains validate HyperEvent’s core innovation: modeling interdependencies among correlated events within a hyper-event significantly enhances dynamic link prediction. HyperEvent uniquely captures holistic structural-temporal patterns that node-centric or isolated-event embeddings fail to represent, particularly in high-churn networks with complex event clusters.
1) High Temporal Surprise: With a surprise metric of 0.987 (the highest among datasets), Review exhibits highly unpredictable interactions. HyperEvent’s reliance on historical event correlations becomes less effective when future events deviate substantially from past structures.
2) Sparse Event Context: Predicting “which product a user reviews next” relies heavily on individual user behavior rather than inter-event dependencies. This misalignment with HyperEvent’s strength in modeling group dynamics reduces its relative advantage.
The exceptional performance on Coin, Comment, and Flight—datasets featuring dense event streams, and strong temporal dependencies reinforces HyperEvent’s superiority in scenarios where event interdependencies drive evolutionary patterns. HyperEvent establishes a new paradigm for temporal graph representation learning, particularly in applications demanding holistic event-sequence reasoning.
4.4 Effectiveness of the Training Algorithm
Method Metric
Training Time per Epoch GPU Memory MRR
TNCN(SOTA) 10h 34m 37s 15,189MB 82.0
HyperEvent(ours) 1h 4m 33s 2,346MB 87.7
Improvement 89.83% $\downarrow$ 84.55% $\downarrow$ 6.95% $\uparrow$
Table 2: Efficiency of HyperEvent on Flight dataset.
Figure 5: GPU memory usage and per-epoch training duration of HyperEvent across varying segment counts ($n_{\text{segment}}$).
The efficiency gains of HyperEvent are quantified in Table 2, contrasting HyperEvent with TNCN – the current SOTA approach topping the TGB Flight dataset leaderboards. HyperEvent achieves a remarkable 89.83% reduction in per-epoch training time (1h 4m 33s vs. TNCN’s 10h 34m 37s) and an 84.55% reduction in GPU memory consumption (2,346MB vs. 15,189MB), while simultaneously improving prediction accuracy with a 6.95% higher MRR (87.7 vs. 82.0). These results unequivocally validate the efficacy of HyperEvent’s training algorithm. The efficiency stems from three key innovations: First, segmenting the event stream into parallelizable segmentations reduces the iterations per epoch by a factor of $n_{\text{segment}}$. Second, a tensorized implementation leverages GPU parallelism directly, bypassing multi-threading bottlenecks and CPU-GPU communication latency. Third, the low-dimensional design of the event correlation vectors (only 12 dimensions) and the lightweight classification model significantly reduce the memory footprint.
Collectively, these results demonstrate HyperEvent’s dual breakthrough: it establishes a new SOTA in prediction accuracy while revolutionizing computational efficiency.
4.5 Modulating Long- and Short-Term Preferences via $n_{\text{neighbor}}$
This divergent behavior originates from the role of $n_{\text{neighbor}}$ in defining the temporal scope for constructing hyper-events, as introduced in Section 3.3. Here, $n_{\text{neighbor}}$ represents the number of most recent interaction partners stored chronologically in the adjacency table, which directly determines the historical length considered for each node. In HyperEvent, this parameter modulates the input to the correlation-based event association calculations from equations (9) to (16), thereby balancing long-term preferences, such as persistent node affinities over extended timelines, against short-term preferences that focus on immediate, transient interactions. Datasets like Coin and Flight, characterized by slowly evolving patterns, achieve higher MRR with larger $n_{\text{neighbor}}$ values, as they leverage richer historical context to accurately identify true hyper-events. In contrast, datasets such as Review and Comment, which involve rapid, high-frequency interactions, perform better with shorter histories, as excessive context may introduce noise that obscures recent relevant events. The findings underscore the adaptability of HyperEvent in capturing dynamic graph dynamics, highlighting how tuning $n_{\text{neighbor}}$ effectively addresses the inherent temporal heterogeneities across applications.
Acknowledging space limitations, comprehensive ablation studies investigating the influence of $n_{\text{latest}}$, the contribution of the Enhanced Correlation Vector, and the impact of the Efficient Training Algorithm—are deferred to the supplementary material.
5 Conclusion
Dynamic link prediction in CTDGs faces critical limitations: existing methods neglect intrinsic correlations within composite events and suffer from scalability constraints for large-scale event streams. This paper introduces HyperEvent, a novel framework that fundamentally reframes the prediction task as hyper-event recognition. By defining hyper-events as cohesive sets of correlated temporal events, HyperEvent shifts the paradigm from isolated link prediction to identifying holistic event structures. Our approach employs efficiently computable event correlation vectors to construct interpretable association sequences, explicitly capturing relational patterns between historical events and the query event. Furthermore, we develop a highly scalable parallel training algorithm leveraging precomputed adjacency tables and event stream segmentation, enabling efficient processing of massive graphs. Extensive experiments demonstrate that HyperEvent achieves competitive prediction accuracy while significantly accelerating training throughput.
Appendix A Complexity Analysis
This section provides a detailed analysis of the computational efficiency of the HyperEvent framework, examining both time and space complexity characteristics. The analysis reveals significant advantages over existing temporal graph learning approaches, particularly in enabling large-scale processing.
Time Complexity per query event prediction is decomposed as follows: i) Adjacency Table Maintenance: Requires $\mathcal{O}(1)$ operations per event insertion due to fixed-length neighbor storage, enabling real-time updates. ii) Hyper-event Extraction: Achieves $\mathcal{O}(1)$ lookup by directly accessing precomputed neighbor lists in $\mathcal{A}$, eliminating neighborhood sampling costs. iii) Relational Vector Calculation: Consumes $\mathcal{O}(n_{\text{neighbor}}^{2})$ operations per event due to pairwise neighbor comparisons, but remains independent of overall graph size. With constrained $n_{\text{neighbor}}$ and optimized 2-hop calculation using $\lfloor\sqrt{n_{\text{neighbor}}}\rfloor$ neighbors, this reduces to practical constant time. iv) Discriminator Inference: Transformer processing exhibits $\mathcal{O}((n_{\text{latest}})^{2}\cdot d_{\text{model}})$ complexity per hyper-event, where $n_{\text{latest}}$ is small and the number of model channels $d_{\text{model}}$ is fixed.
The aggregate $\mathcal{O}(1)$ neighborhood operations demonstrate independence from global graph scale $|\mathcal{V}|$ or $|\mathcal{E}|$, enabling constant-time prediction scaling.
Space Complexity is dominated by: i) Adjacency Tables: Require $\mathcal{O}(|\mathcal{V}|\cdot n_{\text{neighbor}})$ storage, linear in node count but with small constant $n_{\text{neighbor}}$ (typically $\leq 50$). ii) Segment Precomputation: Efficient training adds $\mathcal{O}(n_{\text{segment}}\cdot|\mathcal{V}|\cdot n_{\text{neighbor}})$ storage for boundary adjacency tables, trading marginal space overhead for massive parallelism. iii) Discriminator Parameters: Occupy fixed $\mathcal{O}(d_{\text{model}}^{2})$ space independent of graph dynamics.
The $\mathcal{O}(|\mathcal{V}|)$ spatial scaling is substantially efficient, particularly given the multiplicative constant $n_{\text{neighbor}}$ is empirically shown to be small. When combined with the segment-level parallelism, HyperEvent achieves unprecedented scalability: The framework efficiently processes temporal graphs of tens of millions of edges through parallel training, while maintaining constant-time prediction latency for individual queries, overcoming critical bottlenecks in streaming graph learning.
Appendix B Experimental Setup
B.1 Datasets
@@
Table 3: TGB dataset statistics

### v2

@@ III-C HyperEvent––A Simple Structural Encoding Baseline
The design philosophy of HyperEvent prioritizes conceptual clarity and computational efficiency over architectural complexity. Its pipeline, depicted in Figure 2, relies on non-learned, handcrafted feature computations followed by a standard transformer classifier, eschewing the complex custom modules often found in recent works. This deliberate simplicity is a core tenet of its utility as a baseline.
The pipeline consists of four components:1) Real-time Adjacency Table Maintenance: Dynamically captures and updates recent neighborhood interactions to preserve the evolving graph structure. 2) Hyper-Event Extraction: Identifies composite event patterns by jointly analyzing neighborhoods of source and destination nodes. 3) Pairwise Relational Encoding: Generates matrices that explicitly model the structural-temporal dependencies between events using the proposed relational vectors. 4) Authenticity Discrimination: Employs a Transformer-based architecture to evaluate the likelihood that an extracted hyper-event represents a legitimate interaction pattern. Subsequent sections detail the design and implementation of each component.
III-C1 Real-time Adjacency Table Maintenance ( in Figure 2)
For efficiency and real-time processing—key attributes of a practical baseline—HyperEvent maintains a dynamic adjacency table $\mathcal{A}$. For each node $v\in\mathcal{V}$, $\mathcal{A}_{v}$ stores the $n_{\text{neighbor}}$ most recent interaction partners ordered chronologically:
@@ III-C1 Real-time Adjacency Table Maintenance ( in Figure 2)
where $t>\tau_{1}>\tau_{2}>\cdots>\tau_{k}$ and $k\leq n_{\text{neighbor}}$.Here, $n_{\text{neighbor}}$ controls the storage overhead of the adjacency table and critically defines the historical depth considered when constructing the context for a query event, enabling efficient aggregation of relevant events. Specifically, a smaller $n_{\text{neighbor}}$ value prioritizes short-term historical interactions, enhancing sensitivity to immediate dynamics, whereas a larger value incorporates long-term contextual patterns, capturing enduring relationships. Practical event prediction scenarios exhibit distinct preferences: for instance, in highly dynamic environments like social networks where interactions evolve rapidly, a lower $n_{\text{neighbor}}$ (e.g., near 10) is preferable to adapt swiftly to recent changes; conversely, in stable relational settings such as collaboration networks, a higher $n_{\text{neighbor}}$ (e.g., near 50) better utilizes extensive historical data to improve accuracy. Empirical verification on practical applications shows that $n_{\text{neighbor}}=$10-50 suffices for most datasets, balancing computational efficiency and contextual richness.
The adjacency table enables $\mathcal{O}(1)$ event retrieval, eliminating costly neighborhood searches. Moreover, the adjacency table can be rapidly obtained since its update consists of $\mathcal{O}(1)$ element addition operations, without requiring any computations, serving as a crucial foundation for the Efficient Training Algorithm introduced in Section III-D.
@@
III-C2 Context Extraction ( in Figure 2)
@@
The context for the query event is constructed as:
@@
Figure 3: Visualization of an example of the event correlation vector computation in the HyperEvent.
III-C3 Relative Structural Encoding ( in Figure 2)
For each event $e=(u,v,t)\in\mathcal{H}_{e^{*}}$, we compute a relative structural feature vector that captures the relationships between the historical event and the query event $e^{*}=(u^{*},v^{*},t^{*})$. This vector is designed based on the intuitive notions of relative neighborhood relationships we identified earlier.
0-hop Correlation: This measures the direct co-occurrence between nodes. It quantifies how often a node appears in the historical interactions of another node:
@@ III-C3 Relative Structural Encoding ( in Figure 2)
This provides a measure of how directly the nodes involved in historical events are related to the query event’s nodes.
1-hop Correlation: This measures the overlap in the immediate neighborhoods of nodes:
$d^{(1)}(a,b)=\frac{1}{|\mathcal{A}_{a}|\cdot|\mathcal{A}_{b}|}\sum_{i=1}^{|\mathcal{A}_{a}|}\sum_{j=1}^{|\mathcal{A}_{b}|}\mathbb{I}\big[\mathcal{A}_{a}(i)=\mathcal{A}_{b}(j)\big],$ (13)
where $\mathbb{I}$ is the indicator function. This captures how the neighborhoods of the query event’s nodes overlap with historical events.
2-hop Correlation: This extends the neighborhood analysis to include the neighbors of neighbors, providing a broader context:
@@
with $\tilde{\mathcal{A}}^{(2)}_{a}=\bigcup_{x\in\tilde{\mathcal{A}}_{a}}\tilde{\mathcal{A}}_{x}$ and $\tilde{\mathcal{A}}_{a}$ being the top-$\lfloor\sqrt{n_{\text{neighbor}}}\rfloor$ neighbors of $a$. This captures broader structural patterns beyond immediate neighborhoods.
Feature Vector Construction: The relative structural feature vector for each historical event is constructed by concatenating these three correlation measures for all relevant node pairs:
[…]
This 12-dimensional vector captures the relative structural relationships between the historical event and the query event. The entire context for the query event is represented as a sequence of these vectors $\mathbf{R}_{\mathcal{H}_{e^{*}}}=[\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{|\mathcal{H}_{e^{*}}|}]$, preserving the chronological order of events.
III-C4 Classifier
To complete the baseline model, we employ a standard Transformer architecture as the classifier. This choice reflects our goal of using a powerful yet generic model, ensuring that the model’s performance is driven by the quality of the input features rather than specialized architectural innovations. The Transformer processes $\mathbf{R}_{\mathcal{H}_{e^{*}}}$ to predict whether the query event is likely to occur:
$D(\mathcal{H}_{e^{*}})\in\mathcal{E}\big)=\sigma\big(\mathbf{w}^{\top}\text{Transformer}(\mathbf{R}_{\mathcal{H}_{e^{*}}})+b\big),$ (16)
where $\sigma$ denotes the sigmoid function. Binary cross-entropy loss is employed to optimize the classifier:
@@ III-C4 Classifier
where $y=1$ if the query event represents a true interaction, and $y=0$ otherwise.
Figure 4: Parallelized training framework. Adjacency tables precomputed at key timestamps ($t_{1},t_{2},t_{3}$) anchor segmented event streams (Event Stream 1, Event Stream 2, Event Stream 3).
III-D Efficient Training Algorithm
The practical utility of a baseline model also depends on its trainability on large-scale datasets. To ensure our method meets this requirement, we propose an efficient training algorithm that addresses the computational challenges of continuous temporal graphs.
@@ III-D Efficient Training Algorithm
The proposed method fundamentally differs from existing methods through its exclusive dependence on adjacency tables rather than learned node states or historical event embeddings. All critical computations—from context extraction to prediction—operate directly on adjacency tables without maintaining incremental node-specific historical information. This design principle enables segmentation-based training: given an adjacency table at any arbitrary starting timestamp, our method can commence training from that precise temporal point without loss of continuity. Consequently, precomputed adjacency tables are only required at segment boundaries, which are efficiently obtainable through cumulative aggregation.
@@
As illustrated in Figure 4, the proposed method partitions the original event stream $\mathcal{G}$ into multiple non-overlapping segments (e.g., Event Stream 1, Event Stream 2 and Event Stream 3). Each segment corresponds to a distinct time interval demarcated by split points ($t_{1}$, $t_{a}$, $t_{b}$). Crucially, adjacency tables are precomputed at the starting timestamp of every segment, capturing the cumulative graph state up to that point. This precomputation mechanism, detailed in Section III-C1, ensures that the cumulative structural information captured at segmentation boundaries is functionally equivalent to that obtained during continuous sequential processing.
The adjacency tables serve as compressed representations of historical interactions preceding each segment. Our method enables fully parallelized training across all segments by decoupling segment computations through precomputed tables.
The number of segments $n_{\text{segment}}$ is a pivotal hyperparameter governing the granularity of the partitioning scheme. A larger $n_{\text{segment}}$ yields finer segmentation, reducing the average segment length at the cost of increased storage overhead for maintaining a larger number of precomputed adjacency tables. Conversely, a smaller $n_{\text{segment}}$ minimizes storage requirements but places more update steps demands on individual segments. This establishes an explicit trade-off between computational time efficiency and spatial overhead: $n_{\text{segment}}$ allows practitioners to tailor the segmentation strategy to specific hardware constraints. In practice, $n_{\text{segment}}$ is typically determined via empirical tuning to balance these opposing resource considerations while ensuring that segment boundaries do not disrupt critical temporal dependencies. The linear growth in adjacency table storage relative to $n_{\text{segment}}$ underscores the need for careful calibration of this parameter.
@@
Tensorized Implementation. The parallel training paradigm fundamentally leverages tensorization rather than conventional multi-threading implementation. Crucially, all core computations within the framework operate directly on tensors, inherently enabling batched parallel execution on hardware accelerators like GPUs. Specifically, the original event stream tensor, shaped as ($|\mathcal{E}|$, 2), is partitioned into contiguous segments, transforming it into a ($n_{\text{segment}}$, $\lceil|\mathcal{E}|/n_{\text{segment}}\rceil$, 2) tensor, where padding with a designated value (-1) is applied to handle event counts not perfectly divisible by the segment count. Computational operations defined for our method naturally extend by introducing an additional segment dimension; no algorithmic redesign is required. This tensor-centric parallelism bypasses the computational overhead intrinsic to CPU multi-threading and eliminates constraints imposed by finite CPU core counts.
TABLE I: TGB Dataset Statistics
@@
III-E Complexity Analysis
This section provides a detailed analysis of the computational efficiency of our method, examining both time and space complexity characteristics. This analysis confirms that our method achieves a highly favorable complexity profile, a necessary trait for a scalable baseline. Its efficiency stands in contrast to many more complex learning models, while its performance, as shown in Section IV, remains highly competitive.
Time Complexity per query event prediction is decomposed as follows: i) Adjacency Table Maintenance: Requires $\mathcal{O}(1)$ operations per event insertion due to fixed-length neighbor storage, enabling real-time updates.ii) Context Extraction: Achieves $\mathcal{O}(1)$ lookup by directly accessing precomputed neighbor lists, eliminating neighborhood sampling costs.iii) Relative Structural Encoding: Consumes $\mathcal{O}(n_{\text{neighbor}}^{2})$ operations per event due to pairwise neighbor comparisons, but remains independent of overall graph size. With constrained $n_{\text{neighbor}}$ and optimized 2-hop calculation using $\lfloor\sqrt{n_{\text{neighbor}}}\rfloor$ neighbors, this reduces to practical constant time.iv) Classifier Inference: Transformer processing exhibits $\mathcal{O}((n_{\text{latest}})^{2}\cdot d_{\text{model}})$ complexity per context, where $n_{\text{latest}}$ is small and the number of model channels $d_{\text{model}}$ is fixed.
The aggregate $\mathcal{O}(1)$ neighborhood operations and $\mathcal{O}(n_{\text{latest}}^{2})$ model inference demonstrate independence from global graph scale $|\mathcal{V}|$ or $|\mathcal{E}|$, enabling constant-time prediction scaling.
Space Complexity is dominated by: i) Adjacency Tables: Require $\mathcal{O}(|\mathcal{V}|\cdot n_{\text{neighbor}})$ storage, linear in node count but with small constant $n_{\text{neighbor}}$ (typically $\leq 50$).ii) Segment Precomputation: Efficient training adds $\mathcal{O}(n_{\text{segment}}\cdot|\mathcal{V}|\cdot n_{\text{neighbor}})$ storage for boundary adjacency tables, trading marginal space overhead for massive parallelism.iii) Classifier Parameters: Occupy fixed $\mathcal{O}(d_{\text{model}}^{2})$ space independent of graph dynamics.
The $\mathcal{O}(|\mathcal{V}|)$ spatial scaling is substantially efficient, particularly given the multiplicative constant $n_{\text{neighbor}}$ is empirically shown to be small. When combined with the segment-level parallelism described in Section III-D, our method achieves unprecedented scalability: The framework efficiently processes temporal graphs of tens of millions of edges through parallel training, while maintaining constant-time prediction latency for individual queries, overcoming critical bottlenecks in streaming graph learning.
IV experiment
This section details the experimental framework, including datasets, baselines, evaluation protocols, and implementation specifics, followed by systematic analyses of comparative performance, ablation studies. The source code for HyperEvent is publicly accessible on GitHub [^2] https://github.com/jianjianGJ/HyperEvent.
IV-A Datasets